one number is 3 more than twice another number. If their sum is 21, find the numbers
Answer by mananth(12270)

let one number be x
the other number will be 2x+3
...
x+2x+3=21
3x+3=21
-3
3x=21-3
3x=18
/3
x=18/3
x= 3
..
the other number = 2x+3= 6+3=9
...
m.ananth@hotmail.ca